If You Have 45 inches of fencing to use.what dimensions would give her the most area.
just to let you know the perimeter equals 45 not the area

Answer by checkley77(12844)

45/4=11.25 would be the length of each of the 4 sides of this square.
11.25*11.25=126.5625 in^2 is the max. area.
Proof:
Try 11.2*11.3=126.56